The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 41 February 13, 2000

Contents

News  1
Abstracts of 13 refereed papers 2
Conference contribution  11
2 job opportunities  12

News

Website on the Small Magellanic Cloud

An amateur astronomer, Andrew Murrell, has put together an impressive website with information on the Small Magellanic Cloud. His web pages include historic notes, astronomical information, and a list of SMC clusters and images that he took with a small telescope. He invested a lot of effort in his pages and took extra care to provide as accurate information as possible. He is now asking that professional astronomers take a look at his pages and let him know about possible mistakes or send him suggestions.

The URL of Murrell’s web pages is

            http://www.geocities.com/CapeCanaveral/Orbit/2142/index.html

He can be reached at hector@dragon.net.au . If you have a moment to spare please take a look. He’ll welcome your comments.

Thanks,

Eva Grebel

Abstracts of Refereed Papers

Cool dust and gas in the Small Magellanic Cloud

S. Stanimirovic^{1,2}, L. Staveley-Smith^{2}, J. M. van der Hulst^{3}, Tj. R. Bontekoe^{4,5}, D. J .M. Kester^{6} and P. A. Jones^{1}

1University of Western Sydney Nepean, P.O. Box 10, Kingswood, NSW 2747, Australia 2Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia 3Kapteyn Astronomical Institute, Postbus 800, NL-9700 AV Groningen, the Netherlands 4Bontekoe Data Consultancy, Herengracht 2 B, 2312 LD Leiden, the Netherlands 5Astronomical Institute ”Anton Pannekoek”, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam 6Laboratorium voor Ruimteonderzoek, SRON, P.O. Box 800, 9700 AV Groningen, the Netherlands

We present high angular resolution far-infrared images of the Small Magellanic Cloud (SMC). The images were reconstructed from data obtained with the Infrared Astronomical Satellite (IRAS), using the Pyramid Maximum Entropy algorithm. The angular resolutions of the images are 1 arcmin at 12, 25 and 60 μm and 1.7 arcmin at 100 μm, which provides an excellent complement to the recent neutral hydrogen (HI) survey from the Australia Telescope Compact Array which has a resolution of 1.6 arcmin. We use the 60 and 100 μm data, together with a standard dust model, to derive a temperature map and a dust column density distribution. Dust temperatures range from 23 to 45 K which is, on average, warmer than for the Galaxy, though there is a deficit of warm, very small dust grains in the SMC. The total amount of cool dust in the SMC is 1.8_{−0.2}^{+1.3} × 10^{4} M_{⊙}. There is a significant variation of the observed dust-to-gas ratio across the SMC. This may be due to corresponding variations in metallicity, but this explanation is not supported by the limited available data in the literature. Alternatively, there may be a large amount of molecular hydrogen (H_{2}) in the SMC, ∼ 8 × 10^{8} M_{⊙}, and/or a large amount (∼ 10^{6} M_{⊙}) of cold dust (≤ 15 K) not seen by IRAS or COBE/DIRBE. The average dust-to-atomic gas ratio is, by mass, 8.2 × 10^{−5} which is a factor of ∼ 30 below the Galactic value. The spatial power spectrum of the dust column density image can be approximated by P_{d}(k) ∝ k^{−3.1} which shows that, as with the HI power spectrum, there is no preferred scale size for dust clouds. A spatial coherence analysis shows that, over a wide range of spatial scales (∼ > 50 pc), HI and dust are well correlated (> 60 %).

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact sstanimi@naic.edu

Hubble Space Telescope Observations of Oxygen-rich Supernova Remnants in the Magellanic Clouds. II. Elemental Abundances in N132D and 1E0102.2–7219

William P. Blair^{1}, Jon A. Morse^{2}, John C. Raymond^{3}, Robert P. Kirshner^{3}, John P. Hughes^{4}, Michael A. Dopita^{5}, Ralph S. Sutherland^{5}, Knox S. Long^{6}, and P. Frank Winkler^{7}

1 Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 2 Center for Astrophysics and Space Astronomy, Department of Astrophysical and Planetary Sciences, University of Colorado, Campus Box 389, Boulder, CO 80309, USA 3 Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, USA 4 Rutgers University, Department of Physics & Astronomy, 136 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA 5 Mt. Stromlo and Siding Spring Observatories, Australian National University, Private Bag, Weston Creek Post Office, ACT 2611, Australia 6 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA 7 Department of Physics, Middlebury College, Middlebury, VT 05753, USA

We present Hubble Space Telescope (HST) Wide Field and Planetary Camera 2 images and Faint Object Spectrograph data of two young supernova remnants in the Magellanic Clouds, N132D (LMC) and 1E0102.2–7219 (SMC). The spectra cover essentially the entire UV/optical range available to HST, and provide the first true comparison of UV/optical line intensities from astrophysical shocks that do not depend on scalings from different aperture sizes or instruments. For the spectra, we isolated specific knots and filaments that contain fast-moving debris of nuclear-processed material that are devoid of hydrogen and appear to have originated from the cores of the progenitor stars. In N132D we also observed a knot on the outer rim of the remnant that represents a shocked interstellar cloud. In the debris from both remnants, we identify only the elements O, Ne, C, and Mg. We find no evidence for oxygen-burning products, such as S, Ca, Ar, etc., which are seen in Cas A and are expected from models of type II supernovae. We suggest that the progenitor stars of N132D and 1E0102.2–7219 had large, oxygen-rich mantles (perhaps Wolf-Rayet stars), and may be the products of type Ib supernovae. Shock modelling demonstrates systematic differences in the relative abundances in the O-rich debris, possibly pointing to different progenitor masses for these two objects. The shocked interstellar knot in N132D shows that we are probably seeing a range of conditions within the ∼ 1′′ aperture, and that no evidence is present for enrichment by a precursor star wind.

Accepted by: The Astrophysical Journal, Part 1 (July 1, 2000 issue) For preprints, contact wpb@pha.jhu.edu

Also available from the URL http://violet.pha.jhu.edu/~wpb/papers/hstorich/

The MACHO Project 9 Million Star Color-Magnitude Diagram of the Large Magellanic Cloud

C. Alcock, R.A. Allsman, D.R. Alves, T.S. Axelrod, A. Basu, A.C. Becker, D.P. Bennett, K.H. Cook, A.J. Drake, K.C. Freeman, M. Geha, K. Griest, L. King, M.J. Lehner, S.L. Marshall, D. Minniti, C. Nelson, B.A. Peterson, P. Popowski, M.R. Pratt, P.J. Quinn, C.W. Stubbs, W. Sutherland, A.B. Tomaney, T. Vandehei, D.L. Welch (The MACHO Collaboration^{1})

1 affiliations include: LLNL, STScI, ANU, MSSSO, UW, UCSD, UCB, Notre Dame, Sheffield, ESO, Oxford, McMaster

We present a 9 million star color-magnitude diagram (9M CMD) of the Large Magellanic Cloud (LMC) bar. The 9M CMD reveals a complex superposition of different age and metallicity stellar populations, with important stellar evolutionary phases occurring over three orders of magnitude in number density. First, we count the non-variable red and blue supergiants, the associated Cepheid variables, and measure the stellar effective temperatures defining the Cepheid instability strip. Lifetime predictions of stellar evolution theory are tested, with implications for the origin of low-luminosity Cepheids. The highly-evolved asymptotic giant branch (AGB) stars in the 9M CMD have a bimodal distribution in brightness, which we interpret as discrete old populations ( ∼^{>} 1 Gyr). The faint AGB sequence may be metal-poor and very old. Comparing the mean properties of giant branch and horizontal branch (HB) stars in the 9M CMD to those of clusters, we identify NGC 411 and M3 as templates for the admixture of old stellar populations in the bar. However, there are several indications that the old and metal-poor field population has a red HB morphology: the RR Lyrae variables lie preferentially on the red edge of the instability strip, the AGB-bump is very red, and the ratio of AGB-bump stars to RR Lyraes is quite large. If the HB second parameter is age, the old and metal-poor field population in the bar likely formed after the oldest LMC clusters. Lifetime predictions of stellar evolution theory lead us to associate a significant fraction of the ∼1 million red HB clump giants in the 9M CMD with the same old and metal-poor population producing the RR Lyraes and the AGB-bump. In this case, compared to the age-dependent luminosity predictions of stellar evolution theory, the red HB clump is too bright relative to the RR Lyraes and AGB-bump. Last, we show that the surface density profile of RR Lyraes is fit by an exponential, favoring a disk-like rather than spheroidal distribution. We conclude that the age of the LMC disk is probably similar to the age of the Galactic disk.

Accepted by: The Astronomical Journal ^{1} For preprints, contact alves@stsci.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0001435

1excerpted from third and final chapter of D.R. Alves’ Ph.D. dissertation; see also Chap. 1 (Alcock et al. 1999, PASP, 111, 1539)

and Chap. 2 (Alves & Sarajedini 1999, ApJ, 511, 225)

The Progenitor Masses of Wolf-Rayet Stars and Luminous Blue Variables Determined from Cluster Turn-offs. I. Results from 19 OB Associations in the Magellanic Clouds

Philip Massey^{1}, Elizabeth Waterhouse^{2}, and Kathleen DeGioia-Eastwood^{3}

1 KPNO/NOAO, P.O. 26732, Tucson, AZ 85726-6732, USA 2 REU Participant, Summer 1998 3 Dept Physics and Astronomy, Northern Arizona Univ., P.O. 6010, Flagstaff, AZ 86011-6010, USA

We combine new CCD UBV photometry and spectroscopy with that from the literature to investigate 19 Magellanic Cloud OB associations that contain Wolf-Rayet (WR) and other types of evolved massive stars. Our spectroscopy reveals a wealth of newly identified interesting objects, including early O-type supergiants, a high mass double-lined binary in the SMC, and, in the LMC, a newly confirmed LBV (R 85), a newly discovered WR star (Sk−69^{∘}194), and a newly found luminous B[e] star (LH85-10). We use these data to provide precise reddening determinations and construct physical H-R diagrams for the associations. We find that about half of the associations may be highly coeval, with the massive stars having formed over a short period (Δτ < 1 Myr). The (initial) masses of the highest mass unevolved stars in the coeval clusters may be used to estimate the masses of the progenitors of WR and other evolved stars found in these clusters. Similarly the bolometric luminosities of the highest mass unevolved stars can be used to determine the bolometric corrections for the evolved stars, providing a valuable observational basis for evaluating recent models of these complicated atmospheres. What we find is the following: (1) Although their numbers are small, it appears that the WRs in the SMC come from only the highest mass (> 70M_{⊙}) stars. This is in accord with our expectations that at low metallicities only the most massive and luminous stars will have sufficient mass-loss to become WRs. (2) In the LMC, the early-type WN stars (WNEs) occur in clusters clusters whose turn-off masses range from 30M_{⊙} to 100 M_{⊙} or more. This suggests that possibly all stars with mass > 30M_{⊙} pass through an WNE stage at LMC metallicities. (3) The one WC star in the SMC is found in a cluster with a turn-off mass of 70M_{⊙}, the same as for the SMC WNs. In the LMC, the WCs are found in clusters with turn-off masses of 45M_{⊙} or higher, similar to what is found for the LMC WNs. Thus we conclude that WC stars come from essentially the same mass range as do the WNs, and indeed are often found in the same clusters. This has important implications for interpreting the relationship between metallicity and the WC/WN ratio found in Local Group galaxies, which we discuss. (3) The LBVs in our sample come from very high mass stars (> 85M_{⊙}), similar to what is known for the Galactic LBV η Car, suggesting that only the most massive stars go through an LBV phase. Recently, Ofpe/WN9 stars have been implicated as LBVs after one such star underwent an LBV-like outburst. However, our study includes two Ofpe/WN9 stars, BE 381 and Br 18, which we find in clusters with much lower turn-off masses (25 − 35M_{⊙}). We suggest that Ofpe/WN9 stars are unrelated to “true” LBVs: not all “LBV-like outbursts” may have the same cause. Similarly, the B[e] stars have sometimes been described as LBV-like. Yet, the two stars in our sample appear to come from a large mass range (> 30 − 60M_{⊙}). This is consistent with other studies suggesting that B[e] stars cover a large range in bolometric luminosities. (4) The bolometric corrections of early WN and WC stars are found to be extreme, with an average BC(WNE)=−6.0 mag, and an average BC(WC4)=−5.5 mag. These values are considerably more negative than those of even the hottest O-type stars. However, similar values have been found for WNE stars by applying Hillier’s “standard model” for WR atmospheres. We find more modest BCs for the Ofpe/WN9 stars (BC=−2 to −4 mag), also consistent with recent analysis done with the standard model. Extension of these studies to the Galactic clusters will provide insight into how massive stars evolve at different metallicities.

Accepted by: The Astronomical Journal For preprints, contact massey@noao.edu

Also available from the URL ftp://gemini.tuc.noao.edu/pub/massey/mcwr.ps.Z

Young Stellar Populations Around SN 1987A

N. Panagia^{1}, M. Romaniello^{2}, S. Scuderi^{3} and R.P. Kirshner^{4}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 ESO, Karl-Schwarzschild-Straße 2, D-85748 Garching bei München, Germany 3 Osservatorio Astrofisico di Catania, Viale A. Doria 6, I-95125, Catania, Italy 4 Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA

We present the first results of a study of the stellar population in a region of 30 pc radius around SN 1987A, based on an analysis of multi-band HST-WFPC2 images.

The effective temperature, radius and, possibly, reddening of each star were determined by fitting the measured broad band magnitudes to the ones calculated with model atmospheres. In particular, we have determined effective temperatures and bolometric luminosities for 21,995 stars, and for a sub-sample of 2,510 stars we also determined individual reddening corrections. In addition, we have identified all stars with Hα equivalent widths in excess of 8 Å amounting to a total of 492 stars.

An inspection to the HR diagram reveals the presence of several generations of young stars, with ages between 1 and 150 Myr, superposed on a much older field population (0.6−6 Gyr). A substantial fraction of young stars have ages around 12 Myr which is the stellar generation coeval to SN 1987A progenitor. The youngest stars in the field appear to be strong-line T Tauri stars, identified on the basis of their conspicuous (W_{eq} > 8Å) Hα excesses. This constitutes the first positive detection of low mass (about 1–2 M_{⊙}) Pre-Main-Sequence (PMS) stars outside the Milky Way. Their positions in the HR diagram appear to require that star formation in the LMC occurs with accretion rates about 10 times higher than in the Milky Way, i.e., ∼ 10^{−4}M_{⊙}yr^{−1}.

SN 1987A appears to belong to a loose, young cluster 12±2 Myrs old, in which the slope of the present mass function is almost identical to Salpeter’s, i.e., Γ = dlogN/dlogM ≃ −1.25 for masses above 3M_{⊙}, but becomes much flatter for lower masses, i.e., Γ ≃ −0.5.

On a large scale, we find that the spatial distribution of massive stars and low-mass PMS stars are conclusively different, indicating that different star formation processes operate for high and low mass stars. This results casts doubts on the validity of an Initial Mass Function (IMF) concept on a small scale (say, less than 10pc). Moreover, it appears that a determination of the low-mass end IMF in the LMC requires an explicit identification of PMS stars. A preliminary analysis, done for the whole field as a single entity, shows that the IMF slope for the young population present over the entire region is steeper than Γ ≃ −1.7.

Accepted by: The Astrophysical Journal For preprints, contact mromanie@eso.org

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0001476

OGLE Cepheids have Lower Amplitudes in SMC than in LMC

B. Paczynski and B. Pindor

Princeton University Observatory, Peyton Hall, Princeton NJ 08544-1001, USA

We selected cepheids from the OGLE database for the Magellanic Clouds in the period range 10^{1.1} ≤ P ≤ 10^{1.4} days. There were 33 objects in the LMC and 35 in the SMC. We find that the median amplitude of cepheids in the LMC is 18% larger than in the SMC, a 4 sigma effect. This implies that the period - flux amplitude relation is not universal, and cannot be used to measure distances accurately.

Submitted to: The Astrophysical Journal Letters

For preprints, contact bp@astro.princeton.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0001417

Harmonizing the RR Lyrae and Clump Distance Scales — Stretching the Short Distance Scale to Intermediate Ranges?

Piotr Popowski

IGPP/LLNL, L-413, P.O.Box 808, Livermore, CA 94551, USA

I explore the consequences of making the RR Lyrae and clump giant distance scales consistent in the solar neighborhood, Galactic bulge and Large Magellanic Cloud (LMC). I employ two major assumptions: 1) that the absolute magnitude - metallicity, M_{V} (RR) – [Fe/H], relation for RR Lyrae stars is universal, and 2) that absolute I-magnitudes of clump giants, M_{I}(RC), in Baade’s Window can be inferred from the local Hipparcos calibration of clump giants’ magnitudes. A comparison between the solar neighborhood and Baade’s Window sets M_{V} (RR) at [Fe/H] = −1.6 in the range (0.59 ± 0.05, 0.70 ± 0.05), somewhat brighter than the statistical parallax solution. A comparison between Baade’s Window and the LMC sets the M_{I}^{LMC}(RC) in the range (−0.33 ± 0.09, −0.53 ± 0.09). The distance modulus to the LMC is μ^{LMC} ∈ (18.24 ± 0.08, 18.44 ± 0.07). I argue that the currently available information slightly favors the short distance scale but is insufficient to select the correct solutions with high confidence.

Submitted to: Astrophysical Journal Letters

For preprints, contact popowski@igpp.llnl.gov
Also available from the URL http://arXiv.org/ps/astro-ph/0001470

A Study of the Populations of X-ray Sources in the Small Magellanic Cloud with ASCA

Jun Yokogawa^{1}, Kensuke Imanishi^{1}, Masahiro Tsujimoto^{1}, Mamiko Nishiuchi^{1}, Katsuji Koyama^{1,2}, Fumiaki Nagase^{3}, Robin H. D. Corbet^{4}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 CREST, Japan Science and Technology Corporation (JST), 4-1-8 Honmachi, Kawaguchi, Saitama, 332-0012, Japan 3 Institute of Space and Astronautical Science, Yoshinodai, Sagamihara, Kanagawa 229-8510, Japan 4 Code 662, NASA/Goddard Space Flight Center, Greenbelt, MD 20771, USA; Universities Space Research Association

The Advanced Satellite for Cosmology and Astrophysics (ASCA) has made multiple observations of the Small Magellanic Cloud (SMC). X-ray mosaic images in the soft (0.7–2.0 keV) and hard (2.0–7.0 keV) bands are separately constructed, and the latter provides the first hard X-ray view of the SMC. We extract 39 sources from the two-band images with a criterion of S/N > 5, and conduct timing and spectral analyses for all of these sources. Coherent pulsations are detected from 12 X-ray sources; five of which are new discoveries. Most of the 12 X-ray pulsars are found to exhibit long-term flux variabilities, hence they are likely to be X-ray binary pulsars (XBPs). On the other hand, we classify four supernova remnants (SNRs) as thermal SNRs, because their spectra exhibit emission lines from highly ionized atoms. We find that XBPs and thermal SNRs in the SMC can be clearly separated by their hardness ratio (the ratio of the count rate between the hard and soft bands). Using this empirical grouping, we find many XBP candidates in the SMC, although no pulsations have yet been detected from these sources. Possible implications on the star-formation history and evolution of the SMC are presented by a comparison of the source populations in the SMC and our Galaxy.

Accepted by: Astrophysical Journal Supplement

For preprints, contact jun@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

Beppo SAX and Chandra Observations of SAX J0103.2–7209 = 2E 0101.5–7225: A New Persistent 345 s X–ray Pulsar in the SMC

G.L. Israel^{1}, S. Campana^{2}, S. Covino^{2}, D. Dal Fiume^{3}, T.J. Gaetz^{4}, S. Mereghetti^{5},

T. Oosterbroek^{6}, M. Orlandini^{3}, A.N. Parmar^{6}, D. Ricci^{7} and L. Stella^{1}

1 Osservatorio Astronomico di Roma, Via Frascati 33, I–00040 Monteporzio Catone (Roma), Italy 2 Osservatorio Astronomico di Brera, Via Bianchi 46, I–23807 Merate (Lc), Italy 3 TeSRE, C.N.R., Via Gobetti 101, I–40129 Bologna, Italy 4 Harvard–Smithsonian Center for Astrophysics, MS–70, 60 Garden Street, Cambridge, MA 02138, USA 5 IFCTR, C.N.R., Via Bassini 15, I–20133 Milano, Italy 6 Astrophysics Division, Space Science Department of ESA, ESTEC, P.O. Box 299, 2200 AG Noordwijk, The Netherlands 7 SAX Science Data Center, ASI, Viale Regina Margherita 202, I–00198 Roma, Italy

We report the results of a 1998 July Beppo SAX observation of a field in the Small Magellanic Cloud (SMC) which led to the discovery of ∼ 345 s pulsations in the X-ray flux of SAX J0103.2–7209. The Beppo SAX X-ray spectrum is well fit by an absorbed power–law with photon index ∼ 1.0 plus a black body component with kT = 0.11 keV. The unabsorbed luminosity in the 2–10 keV energy range is ∼ 1.2 × 10^{36} erg s^{−1}. In a very recent Chandra observation the 345 s pulsations are also detected. The available period measurements provide a constant period derivative of −1.7 s yr^{−1} over the last three years making SAX J0103.2–7209 one of the most rapidly spinning-up X-ray pulsars known. The Beppo SAX position (30′′ uncertainty radius) is consistent with that of the Einstein source 2E 0101.5– 7225 and the ROSAT source RX J0103.2–7209. This source was detected at a luminosity level of few 10^{35}–10^{36} erg s^{−1} in all datasets of past X–ray missions since 1979. The ROSAT HRI and Chandra positions are consistent with that of a m_{V} = 14.8 Be spectral type star already proposed as the likely optical counterpart of 2E 0101.5–7225. We briefly report and discuss photometric and spectroscopic data carried out at the ESO telescopes two days before the Beppo SAX observation. We conclude that SAX J0103.2–7209 and 2E 0101.5–7225 are the same source, a relatively young and persistent X-ray pulsar in the SMC.

Accepted by: The Astrophysical Journal Letters

For preprints, contact Gianluca.Israel@oar.mporzio.astro.it
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0001423

Testing Hydrodynamic Models of LMC X-4 with UV and X-ray Spectra

B. Boroson^{1}, T. Kallman^{1}, J.M. Blondin^{2}, and M.P. Owen^{2}

1 Goddard Space Flight Center, Greenbelt, MD 20771, USA 2 North Carolina State University, USA

We compare the predictions of hydrodynamic models of the LMC X-4 X-ray binary system with observations of UV P Cygni lines with the GHRS and STIS spectrographs on the Hubble Space Telescope. The hydrodynamic models determine density and velocity fields of the stellar wind, accretion stream, and accretion disk. We use a Monte Carlo code to determine the UV P Cygni line profiles by simulating the radiative transfer of UV photons that originate on the star and are scattered in the wind. The qualitative orbital variation predicted is similar to that observed, although the model fails to reproduce the strong orbital asymmetry (the observed absorption is strongest for ϕ > 0.5). The model predicts a mid-eclipse X-ray spectrum, due almost entirely to Compton scattering, with a factor 4 less flux than observed with ASCA. We discuss how the model may need to be altered to explain the spectral variability of the system.

Submitted to: The Astrophysical Journal

For preprints, contact bboroson@falafel.gsfc.nasa.gov
Also available from the URL http://hea-www.harvard.edu/~bboroson/HomePage.html

Not Enough Stellar Mass Machos in the Galactic Halo

T. Lasserre^{1}, C. Afonso^{1}, J.N. Albert^{2}, J. Andersen^{6}, R. Ansari^{2}, É. Aubourg^{1}, P. Bareyre^{1,4}, F. Bauer^{1}, J.P. Beaulieu^{3}, G. Blanc^{1}, A. Bouquet^{4}, S. Char^{7}, X. Charlot^{1}, F. Couchot^{2}, C. Coutures^{1}, F. Derue^{2}, R. Ferlet^{3}, J.F. Glicenstein^{1}, B. Goldman^{1}, A. Gould^{8,1}, D. Graff^{8,1}, M. Gros^{1}, J. Haïssinski^{2}, J.C. Hamilton^{4}, D. Hardin^{1}, J. de Kat^{1}, A. Kim^{4}, É. Lesquoy^{1,3}, C. Loup^{3}, C. Magneville^{1}, B. Mansoux^{2}, J.B. Marquette^{3}, É. Maurice^{5}, A. Milsztajn^{1}, M. Moniez^{2}, N. Palanque-Delabrouille^{1}, O. Perdereau^{2}, L. Prévot^{5}, N. Regnault^{2}, J. Rich^{1}, M. Spiro^{1}, A. Vidal-Madjar^{3}, L. Vigroux^{1}, S. Zylberajch^{1} The EROS collaboration

1 CEA, DSM, DAPNIA, Centre d’Études de Saclay, F-91191 Gif-sur-Yvette Cedex, France 2 Laboratoire de l’Accélérateur Linéaire, IN2P3 CNRS, Université de Paris-Sud, F-91405 Orsay Cedex, France 3 Institut d’Astrophysique de Paris, INSU CNRS, 98 bis Boulevard Arago, F-75014 Paris, France 4 Collège de France, Physique Corpusculaire et Cosmologie, IN2P3 CNRS, 11 pl. M. Berthelot, F-75231 Paris, France 5 Observatoire de Marseille, 2 pl. Le Verrier, F-13248 Marseille Cedex 04, France 6 Astronomical Observatory, Copenhagen University, Juliane Maries Vej 30, DK-2100 Copenhagen, Denmark 7 Universidad de la Serena, Facultad de Ciencias, Departamento de Fisica, Casilla 554, La Serena, Chile 8 Departments of Astronomy and Physics, Ohio State University, Columbus, OH 43210, USA

We combine new results from the search for microlensing towards the Large Magellanic Cloud by eros2 (Expérience de Recherche d’Objets Sombres) with limits previously reported by eros1 and eros2 towards both Magellanic Clouds. The derived upper limit on the abundance of stellar mass machos rules out such objects as an important component of the Galactic halo if their mass is smaller than 1M_{⊙}.

Submitted to: Astronomy and Astrophysics Letters

For preprints, contact TLasserre@cea.fr, AMilsztajn@cea.fr
Soon available as http://xxx.lanl.gov/abs/astro-ph/0002253

The Warp of the Galaxy and the Orientation of the LMC Orbit

I. García-Ruiz^{1}, K. Kuijken, J. Dubinski^{2}

1 Kapteyn Astronomical Institute, Postbus 800, 9700 AV Groningen, The Netherlands 2 Department of Astronomy and CITA, University of Toronto, 60 St. George Street, Toronto, Ontario M5S 3H8, Canada

After studying the orientation of a warp generated by a companion satellite, we show that the Galactic Warp would be oriented in a different way if the Magellanic Clouds were its cause. We have treated the problem analytically, and complemented it with numerical N-Body simulations.

Submitted to: Monthly Notices of the Royal Astronomical Society For preprints, contact iruiz@astro.rug.nl

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0002057

Chemical enrichment and star formation in the Milky Way disk. II. Sample description and chromospheric age–metallicity relation

H.J. Rocha-Pinto ^{1}, John Scalo^{2}, W.J. Maciel^{1}, and Chris Flynn^{3}

1 Instituto Astronomico e Geofisico, Av. Miguel Stefano 4200, 04301-904 Sao Paulo SP, Brazil 2 The University of Texas at Austin, Austin, USA 3 Tuorla Observatory, Väisäläntie 20, FI-21500 Pikkiiö, Finland

A chromospheric age distribution of 552 late-type dwarfs is transformed into a star formation history by the application of scale height corrections, stellar evolutionary corrections and volume corrections. We show that the disk of our Galaxy has experienced enhanced episodes of star formation at 0–1 Gyr, 2–5 Gyr and 7–9 Gyr ago, although the reality of the latter burst is still uncertain. The star sample birthsites are distributed over a very large range of distances because of orbital diffusion, and so give an estimate of the global star formation rate. These results are compared with the metalenrichment rate, given by the age–metallicity relation, with the expected epochs of close encounters between our Galaxy and the Magellanic Clouds, and with previous determinations of the star formation history. Simulations are used to examine the age-dependent smearing of the star formation history due to age uncertainties, and the broadening of the recovered features, as well as to measure the probability level that the history derived to be produced by statistical fluctuations of a constant star formation history. We show with more than 98% level of confidence that the the star formation of our Galaxy was not constant.

Submitted to: Astronomy & Astrophysics For preprints, contact Helio Rocha-Pinto helio@iagusp.usp.br

Abstracts of Non-Refereed Papers

TP-AGB evolution with overshoot for low-mass stars as a function of metallicity

Falk Herwig^{1}, Thomas Blöcker^{2}, Thomas Driebe^{2}

1 Universität Potsdam, Institut für Physik, Potsdam, Germany 2 Max-Planck-Institut für Radioastronomie, Bonn, Germany

We give a brief review on the properties of asymptotic giant branch stellar evolution models with overshoot. Then we describe new model calculations with overshoot. Initial masses are ranging from 1 to 3M_{⊙} and metallicities are Z=0.02, 0.01 and 0.001. Third dredge-up occurs efficiently for low masses and carbon stars are formed, with some at core masses as low as 0.58M_{⊙}. After the thermal pulse at which stars become C-rich the luminosities are in the range of the observed C-star luminosity function during the whole interpulse phase and for all C-star models. The dredge-up evolution depends mainly on the core mass at the first thermal pulse and on the metallicity. The Z=0.001 models of the 2 and 3M_{⊙} sequence become C-rich almost instantaneously after the onset of the first thermal pulses. For the 2M_{⊙} case the C/O ratio initially exceeds 4. During following dredge-up episodes the C/O ratio decreases.

To appear in: “The Changes in Abundances in Asymptotic Giant Branch Stars”, ed. F. D’Antona and R. Gallino, Mem. Soc. Astron. Ital.

For preprints, contact fherwig@astro.physik.uni-potsdam.de
Also available from the URL http://www.astro.physik.uni-potsdam.de/~fherwig

Job Opportunities

Postdoctoral Fellowships at the American Museum of Natural History, New York

The American Museum of Natural History in New York, USA, is seeking proposals for innovative research in collaboration with museum faculty in the fields of cataclysmic binaries, stellar populations and globular clusters (M. Shara) and the structure of the interstellar medium and star-forming clouds (M.-M. Mac Low).

Computational resources available at the Museum include 12 and 28 processor SGI Onyx Systems at the Hayden Planetarium as well as a 256 processor Beowulf Cluster.

The Fellowships support independent research in association with the resident staff and are not restricted to recent PhDs. Terms are for up to two years. Fellowships include allowances for relocation as well as research and travel support.

Applicants MUST contact the Department of Astrophysics to discuss their interests and background before applying in order to determine departmental interest.

Application on prescribed forms must be received by February 29, 2000.

For information please contact

Dr. Mordecai-Mark Mac Low

Department of Astrophysics

American Museum of Natural History

New York, NY 10024-5192

   E-mail: mordecai@amnh.org

Application forms can also be downloaded from:

   http://research.amnh.org/grants/index.html

Postdoctoral Research Position at the Universidad de Concepción, Chile

The astronomy group of the Universidad de Concepción, Chile is soliciting applications for a postdoctoral position. The successful applicant will work 50% of his/her time with Dr. Wolfgang Gieren on the reduction and analysis of mosaic images of the nearby spiral galaxy NGC 300 to detect and analyze Cepheid variables, which will be used to calibrate the effect of metallicity on Cepheid-based distances to galaxies. A complete set of BVRI wide field images is available for this purpose, which will be complemented by VLT multiobject spectroscopic data on blue supergiants in NGC 300 later this year, which will be used to calibrate the radial metallicity gradient in NGC 300 (in collaboration with R.P. Kudritzki, Munich). The other 50% of the postdoc’s time will be available to conduct independent research.

The successful applicant will have access to all telescopes operating in Chile, including VLT, Gemini and Magellan 1, under the 10% observing time guaranteed to astronomers working in Chilean institutions. There is also a 5% access to Gemini North. The Universidad de Concepción is the most important Chilean university outside Santiago and has recently implemented a strong research program in astronomy (for more details, see http://coma.cfm.udec.cl). Currently, the astronomy group has four staff and one postdoc. Excellent computer facilities will be available to the successful applicant. The position is for two years and can be filled immediately. Extension to a third year might be possible. Salary is US$ 25,000 per year. Applicants must have a strong background in data reduction and analysis using the IRAF package. Previous experience with the DAOPHOT and/or DoPHOT photometry packages would be very advantageous. Applicants should send a curriculum vitae, bibliography, a statement of research interests, and arrange for three reference letters to be sent to:

Dr. Wolfgang Gieren

Astronomy Group

Departamento de Fisica

Universidad de Concepción

Casilla 160-C

Concepción, Chile

   E-mail: wgieren@coma.cfm.udec.cl

Applicant review will begin March 1, 2000, and continue until the position is filled.